# On the Girth of Graph Lifts

### v1

@@ Abstract
The size of the smallest $k$-regular graph of girth $g$ is denoted by the well studied function $n(k,g)$. We suggest generalizing this function to $n(H,g)$, defined as the smallest size girth $g$ graph covering the, possibly non-regular, graph $H$. We prove that the two main combinatorial bounds on $n(k,g)$, the Moore lower bound and the Erdös Sachs upper bound, carry over to the new setting of lifts, even in their non-asymptotic form.
We also consider two other generalizations of $n(k,g)$: i) The smallest size girth $g$ graph sharing a universal cover with $H$. We prove that it is the same as $n(H,g)$ up to a multiplicative constant. ii) The smallest size girth $g$ graph with a prescribed degree distribution. We discuss this known generalization and argue that the new suggested definitions are superior.
We conclude with experimental results for a specific base graph and with some conjectures and open problems.
@@ 1 Introduction
The girth of a graph is the length of its shortest cycle. It is a classical challenge to determine $n(k,g)$, the smallest possible number of vertices $n$ of a $k$-regular graph of girth $g$. Graphs achieving this minimum are called $(k,g)$-cages, see [3, 5] for a survey of the subject. There are two simple combinatorial arguments yielding lower and upper bounds on $n(k,g)$, namely the Moore lower bound and the Erdös Sachs 1963 upper bound, [4]. Their asymptotic version is:
@@
To this day, the Moore bound remains the best asymptotic lower bound, and Erdös Sachs achieves the best asymptotic upper bound by means of a combinatorial construction. There are number theoretic constructions that do better, such as the celebrated 1988 construction by Lubotzky, Phillips and Sarnak [10], that improves the upper bound to $O({(k-1)}^{3g/4})$.
Less is known about the minimal size of non-regular girth $g$ graphs. The main result is a generalization of the Moore bound to non-regular graphs by Alon, Hoory and Linial 2002 [1]: The smallest possible number of vertices of a girth $g$ graph with minimal degree at least two and degree distribution ${\cal D}$ is lower bounded by:
@@
where $\Lambda=\Lambda({\cal D})=\prod_{d}{(d-1)}^{(d\,p_{d})/{(\sum_{d^{\prime}}d^{\prime}p_{d^{\prime}}})}$ and $p_{d}$ is the portion of degree $d$ vertices specified by ${\cal D}$. The base of the exponent, $\Lambda$, is some weighted average of the degree minus one which is known to satisfy $\Lambda\geq\sum_{d}(d-1)p_{d}$.
@@
However, accepting $n({\cal D},g)$ as the generalization of $n(d,g)$ to non-regular graphs has several drawbacks. Firstly, the Erdös Sachs construction does not carry over to the new setting, so there is no ”trivial” upper bound. Secondly, the existence of Moore graphs, achieving equality with the non-asymptotic version of (2), does not seem possible, expect for degenerate cases.
@@
In this work, we argue that there is a better generalization of the problems using lifts of graphs. Given any finite graph $H$ with minimal degree at least two, $n(H,g)$ is the smallest lift of $H$ with girth at least $g$. We prove that the Moore bound and the Erdös Sachs construction naturally generalize to the new setting, even in their non-asymptotic form.
Following is the main theorem of the paper. The required notation will be introduced later. At this point, it suffices to know that $\rho=\rho(H)$ is some average degree minus one of the vertices of $H$. If $H$ is $k$-regular, then $\rho(H)=k-1$.
Theorem 1.
Given a fixed finite graph $H$ where $\rho$ is the growth rate of its universal covering tree:
$\Omega({\rho(H)}^{g/2})\leq n(H,g)\leq O({\rho(H)}^{g})$ (3)
The rest of the paper is organized as follows. In Section 2 we define graph lifts, the universal cover of a graph, its growth rate $\rho$ and we show that $\rho$ is the Perron eigenvalue of non-backtracking adjacency matrix of the graph. In Section 3 we prove the non-asymptotic form of Theorem 1 and derive the Theorem as a corollary. In Section 4 we explore the relation of $n(H,g)$ to similar definitions. Namely, the size of the smallest graph sharing the same universal cover and the smallest graph sharing the same degree distribution. We demonstrate these results for the base graphs $K_{3,2}$ and $H_{23}$, shown in Figure 2 (a) and (b). In Section 5, we give numeric results for upper and lower bounds on $n(H_{23},g)$. Namely, the Moore bound, the Erdös Sachs construction, exhaustive search, and five greedy construction algorithms. We conclude our work with several conjectures and open questions.
2 Lifts of Graphs
In what follows, for the sake of generality, we follow Friedman and others [2, 6, 7, 8] and use a slightly non-standard definition of graphs. Graphs are always undirected and connected, but not necessarily simple, as we allow multiple edges and two types of self loops: half-loops, and whole-loops. Readers that are only interested in simple graphs, can ignore the extra notation and consider each edge as a pair of directed edges, each being the inverse of the other.
A graph $H$ is the tuple $(V,\overline{E},h,t,\iota)$, where $V$ is the vertex set, $\overline{E}$ is the directed edge set, $h,t:\overline{E}\rightarrow V$ respectively map an edge to its head or tail vertex, and the inverse $\iota$, maps an edge to its inverse $\iota(e)=e^{-1}$. The obvious relations hold: $h(e^{-1})=t(e)$ and that the inverse mapping is an involution. An edge $e$ with $h(e)=t(e)$ is called a loop: a half-loop if $e=e^{-1}$ and a whole-loop otherwise. The degree of $v$, denoted $\deg(v)$ is the number of edges whose tail is $v$, i.e. $\deg(v)=|t^{-1}(v)|=|h^{-1}(v)|$. The neighborhood of $v$ is $\{h(e):e\in t^{-1}(v)\}$, namely the heads of edges emanating from $v$. A length $k$ path $P$ from $u$ to $v$ is an edge sequence $e_{1},\ldots,e_{k}$ such that $t(e_{1})=u$, $h(e_{k})=v$ and $h(e_{i})=t(e_{i+1})$ for all $i$. The path $P$ is non-backtracking if $e_{i+1}\neq(e_{i})^{-1}$ for all $i$. Let $\mbox{dist}(u,v)$ be the length of the shortest path from $u$ to $v$. Denote the radius $r$ ball around some vertex $v$ or edge $e$ by $B_{r}(v)=\{u:\mbox{dist}(u,v)\leq r\}$ and $B_{r}(e)=B_{r}(h(e))\cup B_{r}(t(e))$.
Let $H$ be a finite graph. Following the definition of a topological covering we say that the graph $G$ covers $H$, or that $G$ is a lift of $H$, if there is a surjective homomorphism $\pi:G\rightarrow H$ that is a bijection of the edges emanating from $v$ to the edges emanating from $\pi(v)$ for all $v\in V(G)$, i.e. $\pi(t^{-1}(v))=t^{-1}(\pi(v))$. It it not difficult to see that fiber size is fixed for all vertices and edges. Namely, there is an integer $n$, called the height of the lift, such that $|\pi^{-1}(v)|=|\pi^{-1}(e)|=n$ for all $v\in V(G)$ and $e\in\overline{E}$. It is convenient to identify $V(G)$ with $V(H)\times[n]$ and $\overline{E}(G)$ with $\overline{E}(H)\times[n]$, where $[n]=\{1,2,\ldots,n\}$. Furthermore, it is convenient to think of $\pi$ as a mapping $\mbox{perm}_{\pi}$ from $\overline{E}(H)$ to the permutation group $S_{n}$, such that for any $e\in\overline{E}(H)$ and $i\in[n]$ the edge in $\pi^{-1}(e)$ with tail $(t(e),i)$ has its head at $(h(e),\mbox{perm}_{\pi}(e)(i))$. It follows that $\mbox{perm}_{\pi}(e^{-1})=\mbox{perm}_{\pi}(e)^{-1}$, and that $\mbox{perm}_{\pi}$ of a half-loop is an involution.
The universal cover of $H$, denoted by $\tilde{H}$, is the graph covering all lifts of $H$. This graph, which is unique up to isomorphism, is an infinite tree, unless $H$ is a finite tree. One way to construct $\tilde{H}$ is to choose some arbitrary vertex $v_{0}\in V(H)$ and let $V(\tilde{H})$ be the set of all finite non-backtracking walks from $v_{0}$. Two vertices of $\tilde{H}$ are adjacent if one walk extends the other by a single step and the root of the tree is the empty walk $\epsilon$.
We define the growth rate of $\tilde{H}$ by $\rho(\tilde{H})=\lim_{r}|B_{r}(v)|^{1/r}$, where $v$ is some arbitrary vertex in $V(\tilde{H})$. The non-backtracking adjacency operator of $H$ is the 0-1 matrix $B$ of directed edges versus directed edges, where $B_{f,e}=1$ if $h(e)=t(f)$ and $e\neq f^{-1}$. We let $\rho(B)$ denote the largest eigenvalue (the Perron eigenvalue) of $B$.
Claim 2.
For any undirected connected graph $H$, we have $\rho(\tilde{H})=\rho(B_{H})$.
Proof.
(Outline) Let $\widetilde{v_{0}}$ be some fixed vertex in $V(\tilde{H})$, and let $v_{0}=\pi(\widetilde{v_{0}})$. Then the number of vertices at distance $r$ from $\widetilde{v_{0}}$, which is $|B_{r}(\widetilde{v_{0}})\setminus B_{r-1}(\widetilde{v_{0}})|$, is equal to the number of length $r$ non-backtracking walks in $H$ starting at $v_{0}$. Denoting by $\delta:\overline{E}\rightarrow\{0,1\}$ the function that is one if $t(e)=v_{0}$, we get that for all $r>0$:
$|B_{r}(\widetilde{v_{0}})\setminus B_{r-1}(\widetilde{v_{0}})|=||(B_{H})^{r-1}\delta||_{1}.$ (4)
Therefore,
$\rho(\tilde{H})=\lim_{r}|B_{r}(\widetilde{v_{0}})|^{1/r}=\lim_{r}|B_{r}(\widetilde{v_{0}})\setminus B_{r-1}(\widetilde{v_{0}})|^{1/r}=\lim_{r}(||(B_{H})^{r-1}\delta||_{1})^{1/r}=\rho(B_{H}).$
3 Moore and Erdös Sachs Bounds for Graph Lifts
In this section we prove the non-asymptotic form of Theorem 1 by means of Lemma 3 and Corollary 5 for the lower bound, and Lemmata 4 and 6 for the upper bound. The asymptotic form of the Theorem is an immediate consequence.
Lemma 3 (Moore bound for lifts).
For any undirected connected graph $H$, we have:
  • $n(H,g)\geq|B_{\lfloor g/2\rfloor}(v)|$ for all $v\in V(\tilde{H})$ if $g$ is odd
  • $n(H,g)\geq|B_{\lfloor g/2\rfloor}(e)|$ for all $e\in E(\tilde{H})$if $g$ is even
Proof.
Suppose that $g=2r+1$ and that $G$ is a lift of $H$ with girth at least $g$. Then $B_{r}(u)$ is a tree for any vertex $u\in V(G)$. Since the universal covering tree $\tilde{H}$ is also a lift of $G$ with some cover map $\pi$, then, for any $v\in V(\tilde{H})$, we have $|B_{r}(v)|=|B_{r}(\pi(v))|\leq|V(G)|$. The case of $g=2r$ is similar. ∎
Lemma 4 (Erdös Sachs bound for lifts).
@@ 3 Moore and Erdös Sachs Bounds for Graph Lifts
Corollary 5.
@@
Proof of Lemma 4.

### v2

@@ Abstract
The size of the smallest $k$-regular graph of girth at least $g$ is denoted by the well-studied function $n(k,g)$. We introduce an analogous function $n(H,g)$, defined as the smallest size graph of girth at least $g$ that is a lift (or cover) of the, possibly non-regular, graph $H$. We prove that the two main combinatorial bounds on n(k,g)n(k,g)— the Moore lower bound and the Erdös-Sachs upper bound—carry over to the new lift setting.
We also consider two other functions: i) The smallest size graph of girth at least $g$ sharing a universal cover with $H$. We prove that it is the same as $n(H,g)$ up to a multiplicative constant. ii) The smallest size graph of girth least $g$ with a prescribed degree distribution. We discuss this known generalization and argue that the new suggested definitions are superior.
We conclude with experimental results for a specific base graph, followed by conjectures and open problems for future research.
@@ 1 Introduction
The girth of a graph is the length of its shortest cycle. It is a classical challenge to determine $n(k,g)$, the smallest possible number of vertices $n$ of a $k$-regular graph of girth at least $g$. Minimal size $k$-regular graphs of girth $g$ are called $(k,g)$-cages, see [3, 6] for a survey of the subject. There are two simple combinatorial arguments yielding lower and upper bounds on $n(k,g)$, namely the Moore lower bound and the Erdös-Sachs 1963 upper bound, [5]. Their asymptotic version is:
@@
To this day, the Moore bound remains the best asymptotic lower bound, and Erdös-Sachs achieves the best asymptotic upper bound by means of a combinatorial construction. There are number theoretic constructions that do better, such as the celebrated 1988 construction by Lubotzky, Phillips and Sarnak [13], that improves the upper bound to $O({(k-1)}^{3g/4})$.
Less is known about the minimal size of non-regular girth $g$ graphs. The result of Alon, Hoory and Linial 2002 [1] generalizes the Moore bound to non-regular graphs. They proved that the smallest possible number of vertices of a girth $g$ graph with minimal degree at least two and degree distribution $\cal D$} is lower bounded by:
@@
where $\Lambda=\Lambda({\cal D})=\prod_{d}{(d-1)}^{(d\,p_{d})/{(\sum_{d^{\prime}}d^{\prime}p_{d^{\prime}}})}$ and $p_{d}$ is the portion of degree $d$ vertices specified by $\cal D$}. The base of the exponent, $\Lambda$, is some weighted average of the degree minus one, known to be lower bounded by the average degree minus one.
@@
However, accepting $n({\cal D},g)$ as the right analogue of $n(d,g)$ to non-regular graphs has several drawbacks. Firstly, the Erdös-Sachs construction does not carry over to the new setting, so there is no ”trivial” upper bound. Secondly, the existence of Moore graphs, achieving tight equality with the non-asymptotic version of (2), does not seem possible, expect for degenerate cases, as indicated by Eisner and Hoory [4].
@@
In this work, we argue that there is a better way to formulate the problem using lifts of graphs. Given any finite connected graph $H$ with minimal degree at least two, $n(H,g)$ is the smallest lift of $H$ with girth at least $g$. We prove that the Moore bound and the Erdös-Sachs construction naturally generalize to the new setting, even in their non-asymptotic form. For the sake of avoiding complications, we exclude the case of $H$ being a cycle, as it is well understood. We also note that the $n(H,g)$ problem is not a generalization of the $n(k,g)$ problem, since the family of all $k$-regular graphs is not necessarily a lift of some base graph $H$.
@@
Theorem 1.1.
For a finite connected multigraph with loops $H$ with minimal degree at least two and maximal degree greater than two:
$\Omega({\rho(\tilde{H})}^{g/2})\leq n(H,g)\leq O({\rho(\tilde{H})}^{g}),$ (3)
@@
2 Lifts of Graphs and the Growth Rate of Balls
Graphs are always undirected and connected, with minimal degree $\geq 2$ and maximal degree >2>2. The graphs are not necessarily simple, as we allow multiple edges and two types of self loops: half-loops, and whole-loops, following the convention of Friedman and others [2, 7, 8, 11]. If the graph is simple, the extra notation can be ignored, and each edge is regarded as a pair of directed edges, each being the inverse of the other.
A graph $H$ is the tuple $(V,\overline{E},h,t,\iota)$, where $V$ is the vertex set, $\overline{E}$ is the directed edge set, $h,t:\overline{E}\rightarrow V$ respectively map an edge to its head or tail vertex, and the inverse $\iota$, maps an edge to its inverse $\iota(e)=e^{-1}$. The obvious relations hold: $h(e^{-1})=t(e)$ and that the inverse mapping is an involution. Note that while a half-loop contributes one to the degree, a whole-loop contributes two. An edge $e$ with $h(e)=t(e)$ is called a loop: a half-loop if $e=e^{-1}$ and a whole-loop otherwise. The degree of $v$, denoted $\deg(v)$ is the number of edges whose tail is $v$, i.e. $\deg(v)=|t^{-1}(v)|=|h^{-1}(v)|$. The neighborhood of $v$ is $\{h(e):e\in t^{-1}(v)\}$, namely the heads of edges emanating from $v$. A length $k$ walk $\omega$ from $u$ to $v$ is an edge sequence $e_{1},\ldots,e_{k}$ such that $t(e_{1})=u$, $h(e_{k})=v$ and $h(e_{i})=t(e_{i+1})$ for all $i$. The walk $\omega$ is non-backtracking if $e_{i+1}\neq(e_{i})^{-1}$ for all $i$. Let $\mbox{dist}(u,v)$ be the length of the shortest walk from $u$ to $v$ (necessarily, such a walk is always non-backtracking). Denote the radius $r$ ball centered at some vertex $v$ or edge $e$ by $B_{r}(v)=\{u:\mbox{dist}(u,v)\leq r\}$ and $B_{r}(e)=B_{r}(h(e))\cup B_{r}(t(e))$.
Let $H$ be a finite undirected and connected graph, as above. Following the definition of a topological covering we say that the graph $G$ covers $H$, or that $G$ is a lift of $H$, if there is a surjective homomorphism $\pi:G\rightarrow H$ that is a bijection of the edges emanating from $v$ to the edges emanating from $\pi(v)$ for all $v\in V(G)$, i.e. $\pi(t^{-1}(v))=t^{-1}(\pi(v))$. One can verify that fiber size is fixed for all vertices and edges. Namely, there is an integer $n$, which we call the height of the lift, such that $|\pi^{-1}(v)|=|\pi^{-1}(e)|=n$ for all $v\in V(G)$ and $e\in\overline{E}$. We identify $V(G)$ with $V(H)\times[n]$ and $\overline{E}(G)$ with $\overline{E}(H)\times[n]$, where $n]=\{1,2,\ldots,n\}$. Furthermore, we regard $\pi$ as a mapping $\mbox{perm}_{\pi}$ from $\overline{E}(H)$ to the permutation group $S_{n}$, where for any $e\in\overline{E}(H)$ and $i\in[n]$ the edge in $\pi^{-1}(e)$ with tail $(t(e),i)$ has its head at $(h(e),\mbox{perm}_{\pi}(e)(i))$. It follows that $\mbox{perm}_{\pi}(e^{-1})=\mbox{perm}_{\pi}(e)^{-1}$ and that $\mbox{perm}_{\pi}$ of a half-loop is an involution.
The universal cover of $H$, denoted by $\tilde{H}$, is the graph covering all lifts of $H$. This graph, which is unique up to isomorphism, is an infinite tree. One way to construct $\tilde{H}$ is to choose some arbitrary vertex $v_{0}\in V(H)$ and let V(H~)V($\tilde${H}) be the set of all finite non-backtracking walks from $v_{0}$. Two vertices of $\tilde{H}$ are adjacent if one walk extends the other by a single step and the root of the tree is the empty walk $\epsilon$.
We define the growth rate of $\tilde{H}$ by $\rho(\tilde{H})=\lim_{r}|B_{r}(v)|^{1/r}$, where $v$ is some arbitrary vertex in V(H~)V($\tilde${H}). The non-backtracking adjacency operator of $H$ is the 0-1 matrix $B$ of directed edges versus directed edges, where $B_{f,e}=1$ if $t(f)=h(e)$ and $e\neq f^{-1}$. Then by the following propositions, as $H$ has a minimal degree of $\geq 2$ and maximal degree >2>2, the matrix $B$ is non-negative irreducible, so by the Perron-Frobenius Theorem, its spectral radius $\rho(B)$ is an algebraically simple eigenvalue corresponding to strictly positive left and right eigenvectors. Furthermore, all eigenvalues whose absolute value is equal to $\rho(B)$ are algebraically simple.
Proposition 2.1 ([10] proposition 3.3,[4]).
The non-backtracking adjacency matrix $B_{H}$ is irreducible iff $H$ is connected with $\operatorname{mindeg}(H)\geq 2$ and $\operatorname{maxdeg}(H)>2$.
Proposition 2.2 (Perron–Frobenius, Theorem 8.4.4 from [9]).
Given a non-negative irreducible matrix $M\in M_{n}(\mathbb{R})$ for $n\geq 2$, the spectral radius $\rho(M)$ is a positive and algebraically simple eigenvalue. Furthermore, up to scaling there are unique vectors $\mathbf{x},\mathbf{y}$ so that $M\mathbf{x}=\rho(M)\mathbf{x}$ and $\mathbf{y}^{T}M=\rho(M)\mathbf{y}^{T}$; these vectors are strictly positive.
Proposition 2.3 (Corollary 8.4.6 (c) from [9]).
Given a non-negative irreducible matrix $M\in M_{n}(\mathbb{R})$, let $k$ be the number of distinct maximal eigenvalues in absolute value. Then, these eigenvalues are algebraically simple and their values are $\rho(M)e^{2\pi ij/k}$ for $j=0,\ldots,k-1$.
Theorem 2.4.
[…]
$c_{1}\rho(B_{H})^{r}\leq|B_{r}(v)|\leq c_{2}\rho(B_{H})^{r}.$
Corollary 2.5.
For any finite undirected connected graph $H$ with $\operatorname{mindeg}(H)\geq 2$ and $\operatorname{maxdeg}(H)>2$, the growth rate of its universal cover $\tilde{H}$ is equal to the spectral radius of $B_{H}$:
$\rho(\tilde{H})=\rho(B_{H}).$
We provide two distinct proofs for Theorem 2.4. The first utilizes the algebraic structure of the non-backtracking matrix $B_{H}$, relying on its spectral properties as described in Proposition 2.3. The second employs combinatorial counting methods.
[…]
Proof I of Theorem 2.4.
Let $H$ be a graph satisfying the hypotheses, let $\tilde{H}$ be its universal cover, and let $\pi$ be the respective cover map. Let $v$ be an arbitrary vertex in $\tilde{H}$. Then, the size of the radius $r$ ball centered at $v$, is
$|B_{r}(v)|=1+\sum_{i=1}^{r}\Delta_{i},$ (4)
where $\Delta_{i}=\Delta_{i}(v)$ denotes the number of length $i$ non-backtracking walks from $v$. The quantity $\Delta_{i}$ can be expressed using the non-backtracking matrix $B_{H}$ as
$\Delta_{i}=\delta_{v}^{T}(B_{H})^{i-1}\mathbf{1},$ (5)
where $\mathbf{1}$ is the all-one vector indexed by the set of directed edges $\overline{E}$, and $\delta_{v}$ is the indicator vector on the edges emanating from $\pi(v)$, where the entry $(\delta_{v})_{e}$ is one if $t(e)=\pi(v)$ and zero otherwise.
Let $B_{H}=V^{-1}\left[\bigoplus_{j=1}^{m}J_{j}\right]V$ be the Jordan decomposition of $B_{H}$. Since $B_{H}$ is non-negative and irreducible, by Proposition 2.1, it follows by Proposition 2.3 that the maximum-modulus eigenvalues correspond to $k$ scalar $1\times 1$ Jordan blocks. Thus, the decomposition can be rearranged so that the Jordan blocks $J_{p}$ for $p=1,\ldots,k$ are the scalars $J_{p}=(\rho\,\omega_{k}^{p-1})$, where $\rho=\rho(B_{H})$ and $\omega_{k}=e^{2\pi i/k}$ (here $i$ denotes $\sqrt{-1}$).
Moreover, the left and right Perron eigenvectors of $B_{H}$, satisfying $\mathbf{y}^{T}B_{H}=\rho\mathbf{y}^{T}$ and $B_{H}\mathbf{x}=\rho\mathbf{x}$, are the first row of $V$ and first column of $V^{-1}$, i.e. $\mathbf{y}^{T}=(V)_{1,\cdot}$ and $\mathbf{x}=(V^{-1})_{\cdot,1}$.
Denoting the zero padded $J_{j}$ blocks by $\widehat{J_{j}}$, we write $B_{H}$ as
$B_{H}=V^{-1}\left[\sum_{j=1}^{m}\widehat{J_{m}}\right]V=V^{-1}\left[\sum_{p=1}^{k}E_{p}\rho\,\omega_{k}^{p-1}+R\right]V,$
where the matrix $E_{p}$ is all-zero except for a one in its $p$-th diagonal entry, and the residue matrix $R=\sum_{j=p+1}^{m}\widehat{J_{j}}$ satisfies the condition $\rho^{\prime}=\rho(R)<\rho$.
Applying this decomposition to the expression for $\Delta_{i}$ yields:
$\Delta_{i}$ $=\rho^{i-1}\sum_{p=1}^{k}(\delta_{v}^{T}V^{-1}E_{p}V\mathbf{1})\,\omega_{k}^{(p-1)(i-1)}+\delta_{v}^{T}V^{-1}R^{i-1}V\mathbf{1}$
$=\rho^{i-1}\sum_{p=1}^{k}a_{p}\,\omega_{k}^{(p-1)(i-1)}+r_{i},$
where $a_{p}=\delta_{v}^{T}V^{-1}E_{p}V\mathbf{1}$ and the residue term $r_{i}=\delta_{v}^{T}V^{-1}R^{i-1}V\mathbf{1}$.
The residue term $r_{i}$ is bounded by:
$|r_{i}|\leq\|\delta_{v}\|\rho(V^{-1})\rho(R)^{i-1}\rho(V)\|\mathbf{1}\|=\Big(\|\delta_{v}\|\rho(V^{-1})\rho(V)\Big)(\rho^{\prime})^{i}=o(\rho^{i})\,\,\text{ as }i\rightarrow\infty.$
Furthermore, $a_{p}$ are complex constants, and the coefficient $a_{1}$, corresponding to the Perron eigenvalue $\rho$, is positive:
$a_{1}=\delta_{v}^{T}V^{-1}E_{1}V\mathbf{1}=\delta_{v}^{T}(V^{-1}E_{1})(E_{1}V)\mathbf{1}=\delta_{v}^{T}\mathbf{x}\mathbf{y}^{T}\mathbf{1}>0.$
Now, let $i-1\equiv s\pmod{p}$, for $s\in\{0,\ldots,k-1\}$. Then, we can rewrite $\Delta_{i}$ as :
$\Delta_{i}=\rho^{i-1}\sum_{p=1}^{k}a_{p}\,\omega_{k}^{(p-1)s}+r_{i}=\rho^{i-1}\widehat{a_{s}}+r_{i},$
where $\widehat{a_{s}}=\sum_{p=1}^{k}a_{p}\,\omega_{k}^{(p-1)s}$. The sequence $\{\widehat{a_{s}}\}_{s=0,\ldots,k-1}$ is not all zero, because the coefficients $\{a_{p}\}_{p=1\ldots k}$ are not all zero and the matrix $\{\omega_{k}^{(p-1)s}\}_{p,s}$ is non-singular, being a Vandermonde matrix (specifically, the discrete Fourier transform matrix).
Since $|r_{i}|=o(\rho^{i})$, and since there exists at least one value $s_{0}$ for which $\widehat{a_{s_{0}}}\neq 0$, it follows that $\Delta_{i}=\Theta(\rho^{i})$ for all $i$ with $i-1\equiv s_{0}\pmod{p}$. Furthermore, as $\Delta_{i}$ is a monotone non-decreasing function of $i$, it follows that $\Delta_{i}=\Theta(\rho^{i})$ for all $i$.
Finally, as the assumptions on $H$ imply that $\rho>1$, the sum in (4) is asymptotically dominated by its last term, $\Delta_{r}$, implying that $|B_{r}(v)|=\Theta(\rho(B_{H})^{r}$ as $r\rightarrow\infty$, as claimed. ∎
Proof II of Theorem 2.4.
As in the first proof, let $H$ be a graph satisfying the hypotheses, let $\tilde{H}$ be its universal cover, let $\pi$ be the respective cover map and let $v$ be an arbitrary vertex in $\tilde{H}$. The non-backtracking adjacency matrix $B_{H}$ is non-negative irreducible by Proposition 2.1, implying that $\rho=\rho(B_{H})$ is an eigenvalue of $B_{H}$ with strictly positive left eigenvector $\mathbf{y}$, satisfying $\mathbf{y}^{T}B_{H}=\rho\,\mathbf{y}^{T}$. By (4) and (5) we express the size of the radius $r$ ball centered at $v$ by
$|B_{r}(v)|=1+\sum_{i=1}^{r}\mathbf{\delta_{v}}^{T}(B_{H})^{i-1}\mathbf{1},$
where $\mathbf{1}$ is the all-one vector indexed by the set of directed edges $\overline{E}$, and $\mathbf{\delta_{v}}$ is the indicator vector on the edges emanating from $\pi(v)$, where the entry $(\mathbf{\delta_{v}})_{e}$ is one if $t(e)=\pi(v)$ and zero otherwise. Also, as $B_{H}$is non-negative irreducible, all edges in $\overline{E}$ are reachable from $v$, implying that for some integer $r_{0}$ the following vector is strictly positive:
$\mathbf{z}^{T}=\sum_{i=1}^{r_{0}}\mathbf{\delta_{v}}^{T}(B_{H})^{i-1}.$
It follows that there is a constant $c_{1}>0$ such that $c_{1}\mathbf{y}\leq\mathbf{z}$ and therefore for all $r\geq r_{0}$ we can lower-bound the size of the radius $r$ ball centered at $v$ by
$|B_{r}(v)|\geq\sum_{i=r-r_{0}+1}^{r}\mathbf{\delta_{v}}^{T}(B_{H})^{i-1}\mathbf{1}=\mathbf{z}^{T}(B_{H})^{r-r_{0}}\mathbf{1}\geq c_{1}\rho^{r-r_{0}}\mathbf{y}^{T}\mathbf{1}\geq\Omega(\rho^{r}).\\$
As for the upper bound, as $\mathbf{y}$ is positive, there is a constant $c_{2}>0$ so that $\mathbf{\delta_{v}}\leq c_{2}\mathbf{y}$, and therefore
$|B_{r}(v)|\leq 1+c_{2}\sum_{i=1}^{r}\mathbf{y}^{T}(B_{H})^{i-1}\mathbf{1}=1+c_{2}(\mathbf{y}^{T}\mathbf{1})\sum_{i=1}^{r}\rho^{i-1}.$
By the hypotheses on $H$ we have $\rho=\rho(B_{H})>1$, implying that the geometric sum is $O(\rho^{r})$. ∎
3 Moore and Erdös-Sachs Bounds for Graph Lifts
In this section we prove the non-asymptotic form of Theorem 1.1 by means of Lemma 3.1 for the lower bound and Corollary 3.3 for the upper bound. The asymptotic result of Theorem 1.1 follows by combining the non-asymptotic results with Theorem 2.4, stating that the size of the radius $r$ ball in the universal cover grows as $\Theta(\rho^{r})$.
Lemma 3.1 (Moore bound for lifts).
Let $H$ be a finite undirected connected graph of minimal degree at least two. Then:
  • $n(H,g)\geq\max_{v\in V(\tilde{H})}|B_{\lfloor g/2\rfloor}(v)|$ if $g$ is odd
  • $n(H,g)\geq\max_{e\in E(\tilde{H})}|B_{\lfloor g/2\rfloor}(e)|$ if $g$ is even
Proof.
[…]
Lemma 3.2 (Erdös-Sachs bound for lifts).
Let $H$ be a finite undirected connected graph of minimal degree at least two and let $T$ be a spanning tree of $H$. Then for any integer $g\geq 2\mbox{diam}(T)+2$ there exists a graph $G$ covering $H$ with $\mbox{girth}(G)\geq g$ and $\mbox{diam}(G)\leq g+2\mbox{diam}(T)$.
Corollary 3.3.
Let $H$ be a finite undirected connected graph of minimal degree at least two, let $T$ be some spanning tree of $H$. Then for any integer $g\geq 2\mbox{diam}(T)+2$:
$n(H,g)\leq\min_{v\in V(\tilde{H})}|B_{g\,+\,2\mbox{diam}(T)}(v)|.$
Proof of Lemma 3.2.